OpenStudy (anonymous):

solve the system by the substitution method 5x+8y=4 x+y=2

sam (.sam.):

Take equation 2, x+y=2 x=2-y ---------------------- Substitute x=2-y into equation 1, 5x+8y=4 5(2-y)+8y=4 10-5y+8y=4 10+3y=4 3y=-6 y=-2 ----------------------- x=2-y x=2-(-2) x=2+2=4